OpenStudy (anonymous):

A circular labyrinth has a circumference of 75 ft. What is the distance from the outside edge of the labyrinth to it's center? Use 3.14 for pi. Round your answer to the nearest foot?

OpenStudy (ghazi):

circumference = 2*3.14*r 75=2.3.14*r and r= 75/6.28

OpenStudy (ghazi):

11.9426=11.94

OpenStudy (anonymous):

Would 12 be good?

OpenStudy (ghazi):

yea you can say that ...if precession is not that important ..12 would be okay

OpenStudy (anonymous):

okay, thank you :)

OpenStudy (ghazi):

:)